### v1

# Space to Policy: Scalable Brick Kiln Detection and Automatic Compliance Monitoring with Geospatial Data

Abstract.
Air pollution is responsible for approximately 7 million deaths annually. The brick manufacturing industry significantly contributes to both economic development and environmental challenges, accounting for 8–14% of air pollution in India. Brick kilns provide employment to millions of workers, raising concerns related to the United Nations Sustainable Development Goals (UN SDGs), such as Goal 8.7 (modern slavery). Policymakers have implemented various regulations and compliance measures for brick kilns to mitigate air pollution and associated health effects. Emission inventories quantifying the contribution of different sectors, including but not limited to traffic, construction, brick kilns, and household burnings are critical for air quality modeling and source apportionment studies. However, the largely unorganized nature of the brick kiln sector necessitates labor-intensive survey efforts, hindering compliance assessment and emission inventory development. Recent efforts by air quality researchers have relied on manual annotation of brick kilns using satellite imagery to build emission inventories, but this approach lacks scalability. Machine-learning-based object detection methods have shown promise for detecting brick kilns; however, previous studies often rely on costly high-resolution imagery and fail to integrate the analysis with governmental policies. To address these limitations and support evidence-based decision-making, we developed a scalable machine-learning pipeline that detected and classified brick kilns across five states in the Indo-Gangetic Plain using a free-for-research, moderate-resolution satellite imagery. Our study covers an area of over 520,000 km2 with a population of 448 million. We validated our detections with on-ground surveys, achieving a high correlation of 0.94. Additionally, we performed automated compliance analysis based on government policies and found that a considerable number of brick kilns violate at least one compliance rule. We find that, in the Delhi airshed, stricter policy enforcement and heightened focus on air pollution have led to the adoption of more efficient brick kiln technologies. Using chemical transport modeling (WRF-CAMx), we conducted a PM2.5 source apportionment study, estimating the brick kiln contribution to be 8% in the Delhi airshed. Furthermore, over 30 million people live within 800 meters of brick kilns, contravening government policies. Acknowledging the socio-economic importance of the brick kiln sector, this study highlights the need for inclusive policies that balance environmental sustainability with the livelihoods of workers. By leveraging advances in machine learning and satellite imagery, this work provides tools to foster cleaner air, healthier communities, and a more sustainable future.
Scope of Applicability
This study is conducted for academic and research purposes to contribute to scientific understanding and inform policy development on air quality and sustainability issues. While the findings highlight opportunities for improving compliance and technology adoption in the brick kiln sector, they are not intended for direct enforcement or punitive actions against individual entities or groups. We acknowledge the socio-economic importance of brick kilns and recommend that any policy interventions based on this work be developed in consultation with all stakeholders, including kiln operators, workers, policymakers, and researchers.
1. Introduction
Air pollution causes seven million deaths annually with 22% occurring in India (UNEP, 2019). In 2019, 99% of the world’s population lived in areas that did not meet WHO air quality guideline levels (World Health Organization, nd). In India, brick kilns contribute 8–14% of air pollution (WorldBank, 2020) and employ 15 million workers, including children (Rajarathnam et al., 2014; Boyd et al., 2018), directly impacting UN Sustainable Development Goal 8.7 (Boyd et al., 2018).
Initially, many brick kilns operated with Fixed Chimney Bull’s Trench Kiln (FCBK), an older technology. Studies show that converting to Zigzag technology can reduce CO and particulate matter emissions by 70% (Tibrewal et al., 2023). FCBKs and Zigzag kilns account for 75% of the brick production in India (Rajarathnam et al., 2014). The National Clean Air Program (NCAP), launched in 2019, introduced guidelines to mitigate air pollution and identified 130 cities as “non-attainment cities” due to non-compliance with air quality guidelines. Indian central government and state governments have established compliance rules for brick kilns to regulate their operations (Bri, [n. d.]). Source apportionment studies and air quality modeling rely on accurate emission inventory (Guttikunda et al., 2019). Developing emission inventories and monitoring compliance require accurate geo-mapping of brick kilns and their technologies. However, brick kilns are small-scale, unorganized units, making geo-mapping through field surveys resource-intensive.
Air quality researchers often use tools like ‘Google Earth’ to manually annotate brick kilns in satellite imagery for emission inventory development. Manually annotating an area of size 3600 km2 takes around 6-8 hours. Scaling this for a country like India would require more than 7000 hours of manual effort. Periodically updating the inventory would again require a similar effort. Recent studies (Redmon et al., 2016; Szegedy et al., 2017; Boyd et al., 2018; Nazir et al., 2020; Lee et al., 2021; Haque et al., 2022; Imaduddin et al., 2023) have demonstrated the potential of object detection methods to scale brick kiln detection from satellite imagery. In our work, we employ state-of-the-art object detection models to detect brick kilns and identify their technologies. To the best of our knowledge, ours is the first large-scale study that accurately detects brick kilns using freely available moderate-resolution imagery. Ours is also the first study that combines brick kiln detection with compliance monitoring in close alignment with government policies and systematically analyzes the growth of brick kilns over time.
[…]
Using the geo-located brick kiln dataset, we applied various big data resources for automated compliance monitoring, focusing on two types of compliance rules: distance-based and technology-based. Our analysis of distance-based compliance, using OpenStreetMap data (OpenStreetMap contributors, 2024b) across 13 categories, revealed that 70% of brick kilns violate at least one rule. For technology-based compliance, we examined trends in brick kiln technologies over 12 years in two non-attainment city airsheds, Delhi (the capital city of India) and Lucknow (the capital city of Uttar Pradesh) using Esri’s historical high-resolution imagery (esr, [n. d.]). Between 2017 and 2021, brick kilns around Delhi significantly shifted toward advanced technologies, reflecting stricter policy enforcement. In contrast, no such shift was observed in Lucknow, indicating weaker implementation of regulations. We further estimated daily emissions for each state using experimentally established emission rates (Rajarathnam et al., 2014). These emissions, combined with other sources, were incorporated into an emission inventory to perform a source apportionment study for Delhi March and April 2024 using WRF-CAMx (camx, 2019). Our findings indicate that brick kilns contribute approximately 8% to PM2.5 concentrations in Delhi. Additionally, using LandScan population data (Lebakula et al., 2024), we determined that 30.66 million people live within 800 meters of brick kilns, violating central government regulations (Bri, [n. d.]).
Our work demonstrates a scalable approach to accelerating policy validation processes by leveraging advancements in machine learning, object detection, satellite imagery, and publicly available big data. Notably, aside from computational resources for model training, all components of our study are cost-free and free from proprietary licenses. This openness enables researchers and policymakers to build upon our framework or conduct similar studies without financial constraints.
This article is structured as follows: Section 2 provides an overview of brick kilns, their impact on air pollution and alignment with UN SDGs, relevant Indian government policies, and object detection models. Section 8 reviews related studies on object detection from aerial imagery, the application of computer vision for UN SDGs, and satellite-based brick kiln detection. In Section 3, we outline the big data resources used for automated compliance monitoring and population exposure estimation. Section 4 details the process of creating a geo-located brick kiln dataset using object detection on satellite imagery. Section 5 presents results from automated compliance monitoring. Section 6 explores the experiments linking air pollution and health effects associated with brick kilns. Section 7 discusses potential policy improvements, and finally, we address the study’s limitations, propose future work, and conclude.
Reproducibility Statement: our work is fully reproducible and we will release all the code and resource links to reproduce the results in this work upon acceptance.
2. Background
In this section, we provide the background on brick kiln operations and technologies and their association with air pollution, sustainability, and Indian government policies.
2.1. Brick Kilns
A brick kiln is a facility where bricks are produced through shaping, drying, firing, and cooling processes. Brick-making requires a sufficient supply of soil, sand, water, and fuel. Soil used for brick-making requires moderate clay content, typically found in deposits along riverbanks, leading to the establishment of many kilns near rivers. The brick-making process involves four main steps (Organisation. and Programme., 1984):
i)
Shaping: Bricks are molded into the desired shapes, either manually or using machines.
ii)
Drying: The bricks are air-dried to allow shrinkage caused by water loss and to strengthen them before stacking for firing.
iii)
Firing: Bricks are heated in kilns to achieve high mechanical strength and water resistance. This step releases polluting gases and particulates into the air through chimneys.
iv)
Exporting: The finished bricks are removed in batches and transported from the kiln.
2.2. Continuous Fire Brick Kilns
Brick kilns are primarily classified by their firing method into two categories: batch production kilns and continuous fire kilns. Continuous fire kilns are more energy-efficient as they reuse warm air from the combustion zone to dry unfired bricks, and the heat contained in the fired bricks is used to preheat the air. In India, approximately 75% of bricks are produced by continuous fire kilns (Weyant et al., 2014), with FCBK and Zigzag kilns accounting for over 75% of continuous fire kilns (Tibrewal et al., 2023). FCBKs feature a closed circular or oval circuit of bricks stacked in the annular space between the kiln’s outer and inner walls. They operate using natural draught generated by a central chimney but suffer from incomplete combustion, leading to significant pollutant emissions (Organisation. and Programme., 1984; Ltd, 2012). Zigzag kilns, on the other hand, have a rectangular design with a chimney positioned at the center or side. The bricks are arranged to guide the fire along a zigzag path, enabling more efficient combustion and better fuel efficiency. The zigzag kilns can operate under natural draught or draught created by an induced draught fan. A majority of the zigzag kilns in India use induced draught fans. Figure 1 illustrates satellite views of Circular Fixed Chimney Bull’s Trench Kilns (CFCBK), FCBKs, and Zigzag kilns.
2.3. Brick Kilns and Sustainability
2.3.1. Air Pollution
In 2017, India’s total brick production was estimated at billion bricks, consuming yr-1 of energy (Organisation. and Programme., 1984). The primary fuels for brick production are coal and biomass, which release polluting gases and particulates, including particulate matter (PM), SO2, CO, and CO2 (Rajarathnam et al., 2014). Brick kilns contribute an estimated 8-14% of India’s air pollution (WorldBank, 2020). Advanced technologies, such as Zigzag kilns, use improved firing methods, reducing PM and CO emissions by approximately 60–70% (Rajarathnam et al., 2014). Approximately 45% of India’s brick production occurs in five states: Uttar Pradesh, Bihar, West Bengal, Haryana, and Punjab, all located in the Indo-Gangetic plain—a region known for its dense population and severe air pollution (Tibrewal et al., 2023). For this reason, our study focuses on these five states.
2.3.2. United Nations Sustainable Development Goals
Brick production relies heavily on labor for shaping bricks, fueling kilns, and maintaining continuous operations (Organisation. and Programme., 1984). In India, the industry employs approximately 15 million workers, including children (Rajarathnam et al., 2014; Boyd et al., 2018). This raises concerns related to forced labor, modern slavery, and human trafficking, which are central issues addressed by UN Sustainable Development Goal (SDG) 8.7 (unG, [n. d.]). Developing a comprehensive inventory of brick kiln locations can facilitate targeted surveys to identify potential human rights violations, provide assistance to workers, and establish clear guidelines for kiln owners to promote ethical practices. Furthermore, our research intersects with the following SDGs:
SDG 3.4 & 3.9: Addressing premature deaths and illnesses from air pollution.
SDG 9.4: Mitigating CO2 emissions from industries.
SDG 11.6: Monitoring PM2.5 and PM10 emissions in cities.
SDG 12.c: Reducing fossil-fuel subsidies.
SDG 13.2: Tracking greenhouse gas emissions.
2.4. Government Policies
Several government policies outline guidelines for regulating brick kilns. The National Clean Air Program (NCAP) (cpc, [n. d.]), launched in 2019, aims to reduce air pollution levels to comply with National Ambient Air Quality Standards (NAAQS). By 2024, 131 cities were identified as non-attainment cities for failing to meet these standards. NCAP has directed states, the Central Pollution Control Board (CPCB), and State Pollution Control Boards (SPCBs) to expedite the adoption of cleaner technologies, such as Zigzag firing technology, in brick kilns (cpc, [n. d.]). In 2017, brick kilns operating without permits or failing to comply with environmental guidelines were shut down in Delhi and the National Capital Region (NCR) (cpc, [n. d.]).
The Environment (Protection) Amendment Rules, 2022 (Bri, [n. d.]), mandated the conversion of all brick kilns within 10 km of non-attainment cities to cleaner technologies, including Zigzag technology, vertical shaft kilns, or piped natural gas, by 2024. The rules also specified:
Brick kilns must be located at least 800 meters from residential areas and fruit orchards.
A minimum distance of 1 km must be maintained between two brick kilns.
States can set additional siting criteria to regulate brick kilns. However, monitoring compliance with these policies remains a challenge due to the limitations of manual surveys, which are resource-intensive. Automated Geographic Information System (GIS) analysis offers a scalable alternative, enabling near-instantaneous compliance monitoring. By minimizing reliance on time-consuming manual methods, GIS tools can enhance efficiency and support targeted inspections for effective policy enforcement.
2.5. Object Detection Models
Object detection, a core task in machine learning-based computer vision, involves identifying objects in an image by predicting their bounding boxes and classifying them into predefined categories. The two types of bounding boxes commonly used in object detection are:
Axis-Aligned Bounding Boxes (AA): Rectangular boxes with edges parallel to the horizontal and vertical axes.
Oriented Bounding Boxes (OBB): Rectangular boxes aligned with the orientation of the objects in the image.
OBBs offer a more precise representation of object areas, especially for irregularly oriented objects. In the context of brick kilns, OBBs facilitate accurate spatial extent estimation, which is essential for assessing brick production capacity and calculating emission factors for air quality modeling. Therefore, we employ object detection models supporting OBB annotations to improve the accuracy and applicability of our analysis.
3. Big Data Resources
In this section, we describe the national and global datasets utilized in our work for brick kiln detection and automatic compliance monitoring.
3.1. Satellite Imagery
Selecting suitable satellite imagery requires balancing cost and computational efficiency. High-resolution commercial imagery, such as Maxar (30-50 cm) and Airbus products, offer detailed views but is often too expensive for research. Sentinel-2 imagery from the European Space Agency offers free 10-meter resolution with a 5-day revisit cycle, making it a cost-effective alternative. Brick kilns, on average, cover 150 meters in latitude and longitude and thus span approximately 15×15 pixels in Sentinel-2 imagery. Detecting them at such a low resolution is challenging. Therefore, we utilized freely available moderate-resolution imagery (4.77 meters) under a research license, as detailed in subsequent sections.
3.1.1. Planet Labs
We utilize satellite imagery from Planet Labs in this work (Labs, [n. d.]). Planet offers multiple satellite products, including 3-meter resolution daily imagery, 4.77-meter monthly and quarterly mosaics111https://desktop.arcgis.com/en/arcmap/latest/manage-data/raster-and-images/what-is-a-mosaic.htm. Imagery is provided through Planet’s Research and Education Program (Labs, [n. d.]) and can be freely downloaded via the Mosaics API (Pla, [n. d.]). Our study area spans five Indian states, covering over 520,000 km² (15.8% of India) area. The large sized images are divided into overlapping sized crops with a 64-pixel overlap (300 meters), ensuring that brick kilns do not get cut at the image edges. We selected imagery from the first quarter of 2024 to align with the peak operational season and favorable weather conditions for cloud-free imagery. Although Planet imagery offers better resolution than Sentinel-2, it remains insufficient for identifying kiln technologies. To overcome this, we annotated high-resolution imagery and transferred the labels to Planet imagery, as detailed in the next section.
3.1.2. Esri Wayback Imagery
Esri, a leading provider of GIS technology (Esri, 2024), delivers high-resolution basemaps (up to 30 cm) through a Web Map Service (WMS). These timestamped basemaps are typically available for one or two dates per month. We utilized Esri imagery from February 8, 2024, to align with the first-quarter Planet mosaics of 2024. We could not use high-resolution Esri imagery to train the object detection models because the WMS endpoint is designed for on-demand access rather than bulk downloads.
3.2. OpenStreetMap
OpenStreetMap (OSM) (OpenStreetMap contributors, 2024a), a community-driven initiative, offers a vast array of geographic datasets contributed by volunteers globally. These datasets include detailed information on infrastructure and natural features, enabling extensive spatial analysis. Using the Geofabrik platform (OpenStreetMap contributors, 2024b), we extracted geometry data for features such as railway tracks, highways, rivers, schools, and habitation. We utilize this data for distance-based compliance monitoring. Table 1 summarizes the extracted features and provides the filters we use to extract the data.
3.3. Government Data
3.4. Population Data
We used the LandScan global population dataset (Lebakula et al., 2024), produced by Oak Ridge National Laboratory, at a resolution of 0.0083 degrees. We download the 2023 version of the dataset. We utilize this dataset to analyze population exposure due to brick kilns’ air pollution, as detailed in Section 6.3.
| Shapefile | Data | Filters |
| gis_osm_landuse_a_free_1.shp | Habitation | fclass = ‘‘residential’’ |
| Orchards | fclass = ‘‘orchard’’ | |
| Nature reserves | fclass = ‘‘nature_reserve’’ | |
[…]
| Hospitals | type = ‘‘hospital’’ | |
| Religious places | type = ‘‘temple’’ or ‘‘mosque’’ or ‘‘church’’ | |
[…]
| State highways | ref starts with ‘‘SH’’ | |
| District highways | ref starts with ‘‘MDR’’ | |
[…]
4. Brick Kiln Dataset
4.1. Initial Data Preparation
An initial labeled dataset is required to train the object detection models. We select the regions for initial labeling with the following criteria: i) highly polluted; ii) highly populated (highly impacted by brick kilns); iii) high likelihood of kilns (as suggested by an air quality expert); and iv) non-attainment cities.
Delhi Airshed, India: Delhi, India’s capital, is highly populated and consistently ranks among the world’s most polluted cities (IQAir, [n. d.]). While no brick kilns are located within Delhi due to strict regulations, the surrounding region has a dense concentration of kilns. We selected an grid over Delhi with a 0.01-degree resolution, as defined by Guttikunda et al. (Guttikunda et al., 2023). The selected region is illustrated in Figure 7(a).
Lucknow Airshed, Uttar Pradesh, India: Lucknow, the capital of Uttar Pradesh (India’s most populous state), is classified as a non-attainment city under NCAP. A airshed grid with 0.01-degree resolution was chosen for annotation, following the definition by Guttikunda et al. (Guttikunda et al., 2023). The selected region is illustrated in Figure 7(b).
West Bengal Small Airshed, West Bengal: A small airshed near the Haldi River in West Bengal was selected due to the high density of brick kilns in the area. A 639 km2 region surrounding the identified kilns was annotated as illustrated in Figure 7(c).
Ahmedabad Buffer Region: Ahmedabad, a non-attainment city, was studied in collaboration with the Gujarat Pollution Control Board (GPCB). Following the Environment (Protection) Amendment Rules, 2022 (Bri, [n. d.]), we geo-located brick kilns within a 10 km buffer of Ahmedabad city. The kilns and selected region is shown in Figure 7(d).
Across these four regions, we manually scanned over 15,018 km2 area and identified 1,621 brick kilns, completing the annotations in 125 hours. Table 2 provides region-wise details of area, annotation time, and class-wise distribution. The selected regions are visualized in Figure 7 in the Appendix.
| Airshed | Area (km2) | CFCBK | FCBK | Zigzag | Total Kilns | Annotation time (hours) |
| Delhi Airshed (Guttikunda et al., 2023) | 6937 | 2 | 35 | 746 | 783 | 58 |
| Lucknow Airshed (Guttikunda et al., 2023) | 3962 | 26 | 225 | 241 | 492 | 33 |
| West Bengal Small | 639 | 0 | 89 | 110 | 199 | 5 |
| Ahmedabad 10 km buffer | 3480 | 38 | 108 | 1 | 147 | 29 |
| Total | 15018 | 66 | 457 | 1098 | 1621 | 125 |
4.2. Annotation Process
The annotation process entails labeling brick kilns with oriented bounding boxes (OBBs) and assigning a category to represent the kiln’s technology. We developed a custom labeling interface using Leafmap (Wu, 2021), as shown in Figure 2, utilizing the Esri Wayback Satellite Imagery WMS layer. The region of interest was divided into a 1 km2 grid (Figure 2(a)), and each grid cell was manually inspected. Within each cell, we draw OBBs around the identified kilns (Figure 2(b)). After completing the annotations, the labeled data were overlaid onto cropped, geo-referenced Planet images. Since both Esri and Planet imagery are in the Web Mercator projection (EPSG:3857), the OBBs remained undistorted during reprojection from Esri to Planet imagery, as demonstrated in Figure 1.
4.3. Model Selection
We utilize YOLO-OBB models provided by Ultralytics (Jocher et al., 2023) for our experiments. These models deliver performance comparable to state-of-the-art OBB detection frameworks. YOLO’s algorithm enables faster inference by detecting objects in a single forward pass, making it well-suited for scaling predictions across large geographic areas. We evaluate all OBB variants (‘YOLOv8’ and ‘YOLO11’) across different sizes (‘n’: nano, ‘s’: small, ‘m’: medium, ‘l’: large, ‘x’: extra-large). The dataset described in Table 2 is combined and split into stratified subsets with an 80:10:10 ratio for training, validation, and testing. Stratification ensures a proportional representation of all classes in each subset. The training set (‘train’) is used to train the models, while the validation set (‘validation’) guides performance monitoring and early stopping. The test set (‘test’) evaluates final model performance. We measure model accuracy using the ‘Weighted mAP50’ (mean Average Precision at 50% Intersection over Union) metric, which adjusts the standard mAP50 by weighting class-specific average precisions (APs) inversely to their test dataset counts. Table 3 summarizes the performance of various models on the test dataset. Among the evaluated models, yolo11m-obb achieves the highest performance and is therefore used for brick kiln detection in subsequent experiments.
| Model | CFCBK | FCBK | Zigzag | Weighted mAP |
| yolov8l-obb | 0.61 | 0.58 | 0.83 | 0.62 |
| yolov8x-obb | 0.63 | 0.55 | 0.82 | 0.63 |
| yolo11x-obb | 0.66 | 0.57 | 0.80 | 0.66 |
| yolo11l-obb | 0.68 | 0.51 | 0.76 | 0.66 |
| yolov8m-obb | 0.68 | 0.54 | 0.79 | 0.66 |
| yolo11m-obb | 0.73 | 0.61 | 0.83 | 0.71 |
4.4. Iterative Dataset Building Process
To expand our dataset, we applied the model trained on the initial data to detect brick kilns in Uttar Pradesh (UP), the most populous state in India and a significant contributor to brick kiln activity. We employed a ‘hand-validation’ process to ensure the accuracy of the model’s predictions. The steps involved in hand-validation are as follows:
If a bounding box is entirely incorrect (does not correspond to a brick kiln), we mark it as a false detection and discard it.
If a bounding box partially overlaps a brick kiln, we adjust it to align accurately with the kiln.
If the detected brick kiln is misclassified, we correct the category label.
We used an IoU (Intersection over Union) threshold of 0.33 for Non-Maximum Suppression (NMS) to eliminate duplicate detections. This threshold balances the need to suppress overlapping boxes while retaining accurate detections. For classification confidence, we set a minimum threshold of 0.25, to gain more recall at the expense of precision.
Out of 26978 detected kilns in Uttar Pradesh, 15627 were found to be correct, yielding a precision of %. While the model’s precision may appear low, it substantially reduces manual annotation effort. Annotating the entire Uttar Pradesh region manually would require over 2000 hours, assuming an average of 15 seconds per annotation. By contrast, the hand-validation of 26978 detections requires approximately 112 hours, representing a significant efficiency improvement.
In the next iteration, we combine the initial dataset with 15627 newly identified and hand-validated kilns and fine-tune the model using a validation split. The updated model is then applied to detect brick kilns across five states: Uttar Pradesh, Bihar, West Bengal, Haryana, and Punjab. The precision of the model in this iteration is 82% for Uttar Pradesh, 76% for Bihar, 66% for West Bengal, 72% for Haryana, and 71% for Punjab. These precision values are derived by dividing the number of correctly identified kilns by the total number of detections in each state. Uttar Pradesh achieves the highest precision, as the updated training dataset includes a significant number of examples from this region, improving the model’s accuracy there. Figure 3 illustrates sample detections, including bounding boxes around each brick kiln, categorized by kiln type. After completing this iterative process, we develop a comprehensive dataset containing a total of hand-validated brick kilns with oriented bounding boxes. Table 4 presents the state-wise distribution of hand-validated brick kilns, while Figure 4 maps the geographic distribution of each kiln technology. From our analysis, we observe the following patterns:
CFCBKs: Predominantly found in Uttar Pradesh.
FCBKs: Present in all five states but most common in Uttar Pradesh.
Zigzag Kilns: Concentrated near the Delhi-NCR region, Bihar, and the eastern part of West Bengal.
This analysis, enriched with geographic and technological classifications, is valuable for further research and policy planning.
in CFCBK_106.png, CFCBK_12.png, CFCBK_131.png, CFCBK_39.png, CFCBK_5.png, CFCBK_50.png, FCBK_0.png, FCBK_1.png, FCBK_10.png, FCBK_13.png, FCBK_14.png, FCBK_17.png, Zigzag_11.png, Zigzag_16.png, Zigzag_18.png, Zigzag_2.png, Zigzag_20.png, Zigzag_22.png
| State | CFCBK | FCBK | Zigzag | Total |
[…]
4.5. External Validation
In 2023, the Uttar Pradesh Pollution Control Board (UPPCB) conducted a comprehensive field survey and prepared a report for judicial proceedings (Uttar Pradesh Pollution Control Board ([n. d.]), UPPCB). This report provides district-wise counts of operational brick kilns, revealing a total of 19,671 brick kilns in Uttar Pradesh as of 2022. Though it provides the addresses of brick kilns, we do not reverse geo-locate them to avoid false mapping.
Figure 5 presents a district-wise comparison of brick kiln counts derived from the UPPCB survey and our dataset. Notably, our study is the first to validate a machine-learning-derived brick kiln dataset using independently collected survey data. A correlation analysis, shown in Figure 5(c), demonstrates strong alignment between the two datasets, with a Pearson correlation coefficient of 0.94. The error analysis reveals a mean error of 41.8, a median error of 33.0, and a standard deviation of 40.7 across districts. These results underscore the robustness of our approach in estimating brick kiln distributions and counts at a district level.
5. Automatic Compliance Monitoring
In this section, we present compliance rules and results from automatic compliance monitoring based on distance-based criteria.
5.1. Distance-based Compliance Monitoring
The Environment (Protection) Amendment Rules, 2022 (Bri, [n. d.]), outline compliance requirements for brick kiln operations across all states of India. Among these, two key distance-based siting rules are: i) brick kilns must maintain a minimum distance of 800 meters from habitation and fruit orchards; ii) brick kilns must be at least 1 kilometer apart from one another. States are permitted to establish additional or stricter siting criteria. For our analysis, if both national and state-specific criteria exist, we prioritize state-specific rules. We examined the compliance criteria for Uttar Pradesh, Bihar, West Bengal, Haryana, and Punjab. Using the big data resources described in Section 3, we automatically computed distances between points of interest (e.g., habitation, hospitals, other kilns) and brick kilns. A summary of the results is provided in Table 5, with the corresponding distance thresholds detailed in Table 9 in the Appendix. We visualize the locations of violating brick kilns in Figure 8. Based on our analysis, the following patterns emerge:
The most frequent violations occur with the ‘Inter-kiln distance’ policy, indicating that brick kilns are predominantly established in clusters.
The second highest violations are associated with the ‘hospitals’ category, which specifies minimum distances from healthcare facilities.
Uttar Pradesh has the most violations as well as the highest percentage of violations.
Punjab and Haryana exhibit fewer violations, partly due to less stringent state-specific siting criteria for some categories.
The clustering of kilns and proximity to habitation are recurrent issues across all studied states. These findings underscore the need for better policy implementations and the utility of automatic compliance monitoring in identifying policy violations systematically and at scale.
| State | |||||||
| Uttar Pradesh | Bihar | West Bengal | Haryana | Punjab | Total | ||
| Criterion | Inter kiln | 6469 | 2343 | 669 | 1035 | 604 | 11120 |
| Hospital | 6741 | 656 | 1271 | 176 | - | 8844 | |
| Habitation | 5765 | 1471 | 362 | 131 | 327 | 8056 | |
| National highway | 1495 | 579 | 128 | - | - | 2202 | |
| River | - | 1029 | 338 | - | - | 1367 | |
| State highway | 720 | 150 | 96 | - | 40 | 1006 | |
| District highway | 695 | - | - | - | - | 695 | |
| Railway | 358 | 122 | 84 | - | - | 564 | |
| Nature reserve | 328 | - | 21 | 0 | - | 349 | |
| Orchard | 120 | 8 | 4 | 0 | 11 | 143 | |
| Wetland | - | 70 | - | - | - | 70 | |
| School | 19 | 15 | 26 | 4 | - | 64 | |
| Religious places | 6 | - | 3 | - | - | 9 | |
| Non compliant | 13296 | 3997 | 2081 | 1162 | 866 | 21402 | |
| Brick Kiln count | 17335 | 6048 | 3105 | 2079 | 2071 | 30638 | |
| Percentage violations | 77 | 66 | 67 | 56 | 42 | 70 | |
5.2. Technology-based Compliance Monitoring
As per the guidelines issued by the Central Pollution Control Board (CPCB) (Board, 2019), new brick kilns in non-attainment cities are required to adopt induced-draft zigzag technology. To track the technology used in brick kilns over time, we leverage high-resolution historical imagery from Esri, which allows us to identify and classify kiln technologies on a year-by-year basis. This analysis can be approached in the following ways:
Automatic method: We can apply a detection model to identify kilns across different time periods, capturing both active and potentially closed kilns. This method requires fine-tuning our model on past imagery to optimize detection accuracy.
Manual method: Alternatively, we manually backtrack the locations of kilns and document their transition or establishment timestamps. This approach misses kilns that may have existed earlier but are now closed. We acknowledge this limitation and discuss recommendations for improving future methodologies in Section 9.
For this work, we opt for the manual method to accurately determine the year of establishment or conversion for each kiln. We employ a binary search strategy (Lin, 2019) on the kilns detected by our model and hand-validated, ensuring a high level of precision in identifying key transitions over time.
We start by examining imagery from 2016, the midpoint of the 2010–2022 timeframe. If kilns were detected in 2016, we then check 2013 (the midpoint between 2010 and 2016) to refine the timeline. We continue narrowing the search in this manner, progressively focusing on shorter periods until we pinpoint the exact year of kiln establishment or technology transition. If kilns were not detected in the selected year, we adjust the timeframe accordingly, moving forward or backward as needed. Using this approach, we classify the kiln technology for each year, beginning in 2010. The resulting counts of kilns by technology over time are shown in Figure 6. Notably, in the Delhi airshed (Figure 6(a)), we observe a sharp transition from Fixed Chimney Bull’s Kilns (FCBK) to Zigzag kilns around 2016, aligning with the stricter enforcement of policies. In Figure 6(b), we observe that Lucknow airshed, including another non-attainment city, Lucknow, has adopted Zigzag technology kilns much earlier than Delhi, but the growth of FCBK and Zigzag has been similar and declining over time. Due to the policy regulations over time, we do not observe a decline in the counts of FCBKs over time in Lucknow airshed. This analysis, enabled by high-resolution historical satellite imagery and advanced detection algorithms, not only reveals technological shifts in kiln practices across regions but also underscores the significant role of data-driven monitoring in tracking the adoption of cleaner technologies.
6. Air Pollution and Health Effects
6.1. Emissions
In this section, we estimate air pollutant emissions from brick kilns. We derive emission rates (g/kg) from previous studies (Rajarathnam et al., 2014; Weyant et al., 2014), as listed in Table 7. Tibrewal et al. (Tibrewal et al., 2023) provide annual brick production data (kg/year). Brick kilns typically operate for six months each year, from 15th November to 15th May in Bihar and West Bengal, and from 15th December to 15th June in Uttar Pradesh, Punjab, and Haryana, based on expert inputs. We calculate daily production by dividing the annual production by 180 days. We estimate emissions in tonnes per day for each state by combining emission rates (g/kg) for each technology with the mass of bricks produced (kg). We allocate the total production among different technologies based on our detections. We performed the calculations as follows (for Uttar Pradesh):
| CFCBK + FCBK | |||
| Zigzag kilns | |||
| Total kilns | |||
| PM2.5 emissions | |||
The final emission estimates are shown in Table 7. This analysis enables policymakers to assess the state-level contribution of brick kilns to various pollutants. Integrating these estimates with precise brick kiln locations can support the development of fine-grained emission inventories and facilitate detailed source apportionment studies as we discuss in the next section.
| State | Mass | PM2.5 | SO2 | CO | CO2 |
| Uttar Pradesh | 794816.67 | 118.51 | 312.33 | 2219.30 | 122759.72 |
| Bihar | 401661.11 | 45.86 | 100.15 | 741.14 | 50890.09 |
| West Bengal | 321627.78 | 39.56 | 91.86 | 670.36 | 43003.29 |
| Haryana | 124283.33 | 11.89 | 21.54 | 167.01 | 13920.39 |
| Punjab | 99488.89 | 10.27 | 20.34 | 154.14 | 11742.67 |
| Total | 1741877.78 | 226.08 | 546.23 | 3951.95 | 242316.16 |
6.2. Source Apportionment
Emissions can be estimated from various sources, but they do not translate to air pollution without accounting for meteorology and complex physicochemical reactions in the atmosphere. PM2.5 and PM10 are major pollutants linked to respiratory diseases in humans (Thangavel et al., 2022). Using guidance from an air quality expert, we estimated PM2.5 and PM10 concentrations in the Delhi Airshed by employing the Chemical Transport Model (CTM), CAMx (camx, 2019). CAMx utilizes meteorological inputs generated by the Weather Research and Forecasting (WRF) model (Skamarock et al., 2019). We ran the model over the Delhi Airshed on an 80 km × 80 grid, covering the months of March and April 2024.
Our findings reveal that brick kilns contributed 8% of the total PM2.5 in the Delhi Airshed during this period. This analysis provides a foundation for creating decision support systems to assess the impact of transitioning brick kiln technologies, such as converting to cleaner Zigzag technology or alternative fuels, on air quality. Such systems can guide policymakers in designing targeted interventions to reduce emissions and mitigate public health risks.
6.3. Population Exposure
In Section 5 and Table 5, we observed that many brick kilns are found close to human habitation. According to an analysis of the Air Quality Life Index conducted by the Energy Policy Institute at the University of Chicago (EPIC) (Energy Policy Institute at the University of Chicago (2024), EPIC), an average citizen living in the Indo-Gangetic plain is likely to lose seven-and-half years of life. We use the global population data ‘LandScan’ from Oak Ridge National Laboratory (Lebakula et al., 2024) to estimate the population living near the brick kilns. As we show in Table 8, 30.66 million people live within 800 meters of brick kilns, which is a non-complying range as per the central government policy (Bri, [n. d.]).
| State | < 0.8 km | < 2 km | < 5 km |
| Uttar Pradesh | 13.81 | 63.32 | 168.83 |
| Bihar | 9.43 | 44.22 | 98.41 |
| West Bengal | 4.35 | 18.54 | 50.47 |
| Punjab | 1.95 | 10.03 | 25.64 |
| Haryana | 1.12 | 6.34 | 19.36 |
| Total | 30.66 | 142.45 | 362.71 |
7. Discussion
Our study demonstrates the potential of machine learning and satellite imagery to monitor compliance and assess the environmental impact of brick kilns at scale. By identifying violations and estimating contributions to air pollution, our work provides actionable insights for policymakers. However, such insights must be contextualized within the socio-economic realities of the brick kiln industry to ensure that interventions are equitable and sustainable. The brick kiln sector is not just a contributor to air pollution but also a critical source of livelihoods for approximately 15 million workers in India. Many of these workers come from underprivileged backgrounds, with limited alternative employment opportunities. Regulatory interventions, if implemented abruptly, risk adversely affecting these vulnerable populations. For instance, stringent enforcement of compliance rules or sudden shutdowns of non-compliant kilns could lead to loss of employment for thousands of workers.
To address this, we suggest the following changes.
We advocate for gradual implementation timelines and targeted support for kiln owners and workers. This includes skill development programs to enable smoother transitions to advanced kiln technologies like Zigzag kilns, which reduce emissions and improve working conditions for the workers.
In this work, we find that various states have uneven policies defined for the brick kilns. In our opinion, policies must be informed by evidence-based studies like ours, which integrate spatial data, emission estimates, and population exposure. This ensures that interventions are effective without being unduly punitive.
The health and safety of workers in brick kilns are often overlooked in policy discussions. Cleaner technologies can potentially mitigate occupational hazards, such as respiratory illnesses caused by prolonged exposure to particulate matter and harmful gases. Complementary measures, such as access to healthcare and education, are essential for long-term social upliftment of the workers.
Our work provides an opportunity to develop sustainable and inclusive policies for the brick kiln sector. By addressing both environmental and human dimensions, this approach aligns with the goals of cleaner air, improved public health, and social equity.
8. Related Work
We discuss related work in three areas: object detection from aerial imagery, the application of satellite imagery for addressing UN Sustainable Development Goals (SDGs), and prior efforts on brick kiln detection.
8.1. Object Detection from Aerial Imagery
[…]
8.2. Computer Vision on Satellite Imagery for UN SDGs
The United Nations’ 17 SDGs have been a focal point for research since their introduction in 2016 (unG, [n. d.]). Satellite imagery and computer vision techniques have been effectively employed to address various SDGs. For instance, Burke et al. (Burke et al., 2021) applied computer vision to large-scale satellite data for agriculture, economics, and informal settlement detection. Xi et al. (Xi et al., 2023) developed a dataset supporting 25 SDG indicators, such as population, nighttime light, poverty, health, and living environment, to aid urban policymaking and SDG monitoring. Kumar et al. (Ayush et al., 2020) used object detection on satellite imagery to measure poverty, aligning with SDGs by informing resource allocation and livelihood improvements. Yeh et al. (Yeh et al., 2020) predicted economic well-being across 20,000 African villages using multi-spectral imagery and deep learning. We have discussed the intersection of our work with UN SDGs in Section 2.3.2.
8.3. Brick Kiln Detection from Satellite Imagery
Various machine learning methods have been applied to detect brick kilns in South Asia. Nazir et al. (Nazir et al., 2020) used a gated neural network combining classification and detection tasks to identify kilns over a 3,300 km2 area with 1,300 kiln images. Lee et al. (Lee et al., 2021) applied deep learning in Bangladesh, identifying 3,345 kilns. A classical machine learning approach using Sentinel-2 imagery classified kilns in Delhi (Imaduddin et al., 2023). Haque et al. (Haque et al., 2022) analyzed heavy metal contamination in soils near brick kilns in Bangladesh. Boyd et al. (Boyd et al., 2018) employed YOLOv3 with axis-aligned bounding boxes (AABBs) to detect fixed chimney bull’s trench kilns (FCBKs) in the Indo-Gangetic plain. However, most previous studies relied on expensive high-resolution satellite imagery, significantly limiting scalability and reproducibility.
Our study marks the first large-scale effort to geo-locate and classify brick kilns by technology type (CFCBK, FCBK, Zigzag) across five states in the Indo-Gangetic Plain, spanning 520,000 km2. We identified kilns, validated our dataset with field surveys, and assessed the evolution of kiln technologies over time. Unlike prior works, we used freely available, moderate-resolution satellite imagery (4.77 m resolution) under an Education and Research license, ensuring scalability, affordability, and reproducibility.
9. Limitations and Future Work
In this section, we outline the limitations of our study and propose directions for future research.
As we establish in Figure 5, our coverage closely resembles the manual survey conducted by the Uttar Pradesh State Pollution Control Board; however, identifying whether a kiln is operational at a given time remains challenging with RGB imagery. During operation, kilns exhibit elevated temperatures, presenting an opportunity to use thermal sensors from satellites for status detection. Current satellite thermal imagery, with resolutions around 100 meters, lacks the precision needed for such applications. Future advancements in satellite technologies may make operational status detection feasible.
In Section 5.2 and Figure 6, we manually analyzed the growth of brick kilns over time for two non-attainment cities. Extending this effort to the state level is resource-intensive and requires manual input. Fine-tuning our detection models on historical satellite imagery could enable automated assessment of kiln technology evolution over time.
We have demonstrated a reasonably accurate precision and practically useful recall of our studies. However, moderate-resolution of the imagery poses a challenge in identifying kiln technologies accurately. One may utilize a high-resolution imagery to improve the accuracy of technology identification.
10. Conclusion
[…]
Appendix A Initial Dataset Regions
Appendix B Automatic Compliance Monitoring
| Criterion | Uttar Pradesh | Bihar | West Bengal | Haryana | Punjab |
| Inter kiln | 800 | 1000 | 300 | 1000 | 1000 |
| Habitation | 1000 | 800 | 800 | 800 | 500 |
| National highway | 300 | 300 | 200 | - | - |
| River | - | 500 | 200 | - | - |
| State highway | 300 | 200 | 200 | - | 100 |
| District highway | 100 | - | - | - | - |
| Railway | 200 | 200 | 200 | - | - |
| Nature reserve | 5000 | - | 5000 | 1000 | - |
| Orchard | 800 | 800 | 800 | 800 | 800 |
| Wetland | - | 500 | - | - | - |
| School | 1000 | 800 | 1000 | 1000 | - |
| Religious places | 1000 | - | 1000 | - | - |